OpenStudy (anonymous):

A set of data items is normally distributed with a mean of 400 and a standard deviation of 60. Find the data item in this distribution that corresponds to the following z-score: z = 3

jimthompson5910 (jim_thompson5910):

z = (x - mu)/(sigma) 3 = (x - 400)/60 3*60 = x - 400 180 = x - 400 ... ... ... x = ??

OpenStudy (anonymous):

x= 580

jimthompson5910 (jim_thompson5910):

yep